OpenStudy (anonymous):

Part 1: Describe the graphs of the functions f(x) = 2x – 3 and g(x) = –2x – 3. Part 2: Compare and contrast the domain and range of f(x) and g(x)

OpenStudy (matheducatormcg):

have you graphed the functions?

OpenStudy (matheducatormcg):

they are both lines in slope intercept form, y=mx+b where m is your slope and b is your y-intercept. do you know how to graph from here?

OpenStudy (anonymous):

kind of sort of

OpenStudy (matheducatormcg):

both lines have b values of -3, which means we can start graphing by plotting a point at -3 on the y-axis. from that point we count out our rise and run from the slope. \[slope=\frac{rise}{run}\] in f(x), m=2 which can be represented as \[\frac{2}{1}\] which means a positive rise of 2 and a run of 1. so from -3 go up 2, over to right 1 and plot a point. now draw a line through both points.

OpenStudy (matheducatormcg):

we do same for g(x), but the m=-2 in g(x) so we represent -2 as \[\frac{-2}{1}\] which is a rise of negative 2 and a run of 1. So, from -3 go down 2 and over to the right 1 and plot a point. draw a line through these points. these graphs should look like an X

OpenStudy (matheducatormcg):

domain represents all possible x values that have corresponding y values in the function range represents all possible y values that have corresponding x values in the function

OpenStudy (matheducatormcg):

compare and contrast seems really boring here. let me know if you want my take considering \[D:x \epsilon R \] x is an element of the real numbers \[R:y\epsilon R\] y is an element of the real numbers